# DiffTester: Accelerating Unit Test Generation for Diffusion LLMs via Repetitive Pattern

### v2

@@ C.4.3 Generated examples of different models on different language tasks
Figure 14,15,16,17,18,19 are the generated examples of two models, DiffuCoder and Dream, on three language tasks–Python, Java and C++. The code segments highlighted in blue represent the repetitive patterns identified within the unit test cases. Notably, these repetitive patterns are consistently observed in each language, indicating their universality across diverse coding environments. This observation suggests that different dLLMs can effectively leverage these repetitive patterns to accelerate the decoding process. By exploiting the presence of such patterns, the models are able to generate test cases more efficiently. These more motivating examples underscore the potential for repetitive pattern utilization as a general mechanism for boosting decoding efficiency in unit test generation scenarios.
C.5 Limitation
A principal limitation of DiffTester is the absence of empirical investigations into the acceleration of UTG on repository-level codebases, which would be of greater practical significance. This omission is primarily attributable to the constrained maximum context lengths supported by existing dLLM architectures. In contrast, repository-scale UTG tasks, such as those exemplified by TestGenEval (jain2025testgenevalrealworldunit), inherently demand the modeling of substantially longer contexts that exceed these architectural limits.
Another limitation arises from the inherent characteristics of certain contemporary code generation models, such as DreamCoder (dreamcoder). Our method was not applied to these models primarily because their generative behavior tends to produce substantial amounts of natural language output, even when explicitly prompted to focus on code generation. This results in a low proportion of code segments, thereby constraining the potential acceleration benefits achievable by our method.
C.6 Algorithm
@@ 5    $merged\leftarrow$ recursively merged root node of $child_{1}$ and $child_{2}$;
/* merged_list represents list of merged node and its source code lines */
@@ 10       $found\leftarrow True$;
 /* Only those greater than 1 indicate the presence of a repetitive pattern. */
@@
Algorithm 2 Unmasking extra tokens via repetitive pattern

### v3

@@ C.4.3 Generated examples of different models on different language tasks
Figure 15,16,17,18,19,20 are the generated examples of two models, DiffuCoder and Dream, on three language tasks–Python, Java and C++. The code segments highlighted in blue represent the structural patterns identified within the unit test cases. Notably, these structural patterns are consistently observed in each language, indicating their universality across diverse coding environments. This observation suggests that different dLLMs can effectively leverage these structural patterns to accelerate the decoding process. By exploiting the presence of such patterns, the models are able to generate test cases more efficiently. These more motivating examples underscore the potential for structural pattern utilization as a general mechanism for boosting decoding efficiency in unit test generation scenarios.
C.5 Algorithm
@@ 5    $merged\leftarrow$ recursively merged root node of $child_{1}$ and $child_{2}$;
/* merged_list represents a list of merged nodes and their source code lines */
@@ 10       $found\leftarrow True$;
 /* Only those greater than 1 indicate the presence of a structural pattern. */
@@
Algorithm 2 Unmasking extra tokens via structural pattern